f(x) = 6 arcsin square root 1-x


1
Expert's answer
2021-01-01T13:45:50-0500

f(x) = "\frac{6arcsinx}{\sqrt{1-x}}"

Evaluate

Multiply the numerator and denominator by the conjugate

"\frac{6arcsin(x)\sqrt{1-x}}{1-x}"